OpenStudy (anonymous):

In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40. How do I begin to solve this?

terenzreignz (terenzreignz):

I remember you :D So... x-intercept... where the graph touches the x-axis. That's fancy talk. To get the x-intercepts, simply set y = 0, and solve. I know you can do this... you've done it quite a few times last time ;)

OpenStudy (anonymous):

The x-intercepts are where the graph crosses the x-axes: or in other words when \(y=0\). So to begin we allow: \(y=3x^2 + 19x – 40=0 \) upon solving this quadratic equation, we have the x-intercepts of the graph!

terenzreignz (terenzreignz):

What he/she said ^ :) If you're stuck, keep this in mind, and you'll never go wrong with quadratics :) Whenever you're faced with solving a quadratic equation in the form ax² + bx + c = 0 Then remember this formula, and you'll be able to regularly get the solutions for x... \[\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

(3x - 5)(x + 8) = 0

terenzreignz (terenzreignz):

Or you could do that :D So, remember how to proceed from here? At least one of these factors must be zero, so check both cases.

OpenStudy (anonymous):

@beesknees01 Excellent!

OpenStudy (anonymous):

x = 5/3 x=- 8

terenzreignz (terenzreignz):

Brilliant ;)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Y intercepts occurs when X=0 and hence X intercepts occur when y=o Therfore if y=0, it follows that.... 0=3X^2=19X-40 using the quadratic formular.....